# Perturbations and Quasinormal Modes of the Dirac Field in Effective Quantum Gravity

### v1

@@ Abstract
Recently, two models of quantum-corrected Schwarzschild-like black holes were developed within Effective Quantum Gravity, and the spectra of bosonic perturbations have been analyzed in several recent studies. In this work, we investigate the quasinormal modes of massless Dirac field perturbations around these black holes. By employing the higher-order WKB method and time-domain integration, we achieved consistency between the two approaches within their common range of validity. This concordance enables us to distinguish the two quantum-corrected black hole models from each other and from their Schwarzschild limit through their quasinormal spectra.
@@ I Introduction
Quasinormal modes (QNMs) a black hole background describe the characteristic oscillations of perturbations in the spacetime surrounding the black hole Nollert (1999); Kokkotas and Schmidt (1999); Konoplya and Zhidenko (2011). These modes are defined by complex frequencies, where the real part corresponds to the oscillation frequency, and the imaginary part is related to the damping rate due to the emission of radiation. Quasinormal modes corresponding to spacetime perturbations are observed with the help of gravitational interferometers Abbott et al. (2016, 2017, 2019). The study of Dirac QNMs is crucial for understanding the stability of fermionic perturbations and provides insights into the behavior of quantum fields in curved spacetime. In the eikonal limit, the QNMs are closely related to the properties of null geodesics, particularly the orbital frequency and the Lyapunov exponent, which characterize the stability of these geodesics Cardoso et al. (2009); Konoplya and Stuchlík (2017); Konoplya (2023); Bolokhov (2024a). However, Dirac QNMs exhibit unique features, such as spin-orbit coupling effects, that distinguish them from scalar or gravitational perturbations. These modes also play a some role in the context of black hole thermodynamics and quantum gravity, where they can be used to probe the underlying quantum structure of spacetime. Analytical and numerical methods, including the WKB approximation and time-domain integration, are commonly employed to calculate these modes, revealing how different black hole parameters, such as mass, charge, and spin, influence the fermionic QNM spectrum. As a result, quasinormal modes of Dirac fields have been extensively studied for a great number of black hole models and gravitational theories Zinhailo (2019); Bolokhov (2024b); González and Vásquez (2014); Konoplya et al. (2023); Jing and Pan (2005); Konoplya and Zhidenko (2018); Rosa and Dolan (2012); Wongjun et al. (2020); Kanti et al. (2006); Al-Badawi and Kraishan (2023); Chen and Jing (2005); Konoplya et al. (2022); Saleh et al. (2016); Chakrabarti (2009); Varghese and Kuriakose (2011); Sebastian and Kuriakose (2014); Lopez-Ortega (2007); Wu and Zhao (2004); Zhang and Jing (2005); Lopez-Ortega (2010); Wang and Gui (2010); Wang et al. (2017); Konoplya and Zhidenko (2007a); Jing (2004); Al-Badawi et al. (2023); Zhou and Zhu (2004); Konoplya et al. (2020); Lopez-Ortega (2014).
@@
Recently, within the Hamiltonian constraints approach, intriguing models for spherically symmetric quantum-corrected black holes that preserve general covariance have been developed in Zhang et al. (2024). As a result, various physical properties of these black holes, such as optical phenomena and particle motion, have been studied in Liu et al. (2024a, b). The quasinormal modes of boson fields have also been recently investigated in Konoplya and Stashko (2024); Bolokhov (2024c); Malik (2024). In this work, we extend these studies to consider massless Dirac field perturbations and calculate the corresponding quasinormal frequencies using two alternative methods: the higher-order WKB method with Padé approximants and time-domain integration. We will demonstrate that the quasinormal modes of the Dirac field exhibit a noticeable shift when the quantum parameter is introduced.
@@ II Black hole metric and wave-like equation
The metric of the quantum corrected black hole is given by the following line element
@@ III.1 WKB Method
This expression provides the square of the QNM frequencies $\omega_{n}$ in terms of the maximum of the potential $V_{0}$, its second derivative $V_{0}^{\prime\prime}$, and the higher-order WKB corrections $\Lambda_{j}$. The sum of higher-order corrections $\sum_{j=1}^{\infty}\Lambda_{j}$ ensures the accuracy of the QNM frequency calculation, especially when going beyond the leading WKB order. The explicit forms of the corrections have been derived in Iyer and Will (1987) up to the third order, and in Konoplya (2003) and Matyjasek and Opala (2017) up to the sixth and 13th orders respectively. The WKB method has been used for finding quasinormal modes in numerous works (see, for example, Konoplya and Zhidenko (2007b); Balart et al. (2023); Bretón et al. (2017); Yu et al. (2022); Konoplya and Zhidenko (2022); Al-Badawi (2023); Xiong et al. (2022); Kokkotas et al. (2011)), so that we will not discuss it here in more detail. Here we will use the 6th order WKB method Konoplya (2003) with Padé approximants Matyjasek and Opala (2017); Konoplya et al. (2019).
@@ III.2 Time-Domain Integration
$\Psi(N)=\Psi(W)+\Psi(E)-\Psi(S)-\Delta^{2}V(S)\frac{\Psi(W)+\Psi(E)}{4}+\mathcal{O}\left(\Delta^{4}\right),$ (9)
@@ IV Quasinormal modes
Perturbations in a spherically symmetric background can be reduced to the wave-like equation with the effective potential which can be approximated in the following way:
@@
In Cardoso et al. (2009), it was demonstrated that there is a correspondence between the parameters of null geodesics (orbital frequency and Lyapunov exponent) and the real and imaginary parts of quasinormal modes in the eikonal limit. This correspondence is known to hold in the majority of cases. However, there are several exceptions and limitations, as discussed in Konoplya and Stuchlík (2017); Konoplya (2023); Bolokhov (2024a). The eikonal limit is interesting not only because in this regime the frequencies could be found in analytic form (see, for instance, Konoplya and Abdalla (2005); Bolokhov (2024d); Konoplya (2002); Zhidenko (2004); Allahyari et al. (2019)), but also because it may bring a catastrophic instability Konoplya and Zhidenko (2017a); Takahashi (2011); Konoplya and Zhidenko (2017b); Gleiser and Dotti (2005); Dotti and Gleiser (2005) in theories with higher corrections in curvature.
@@
For the present case of Dirac field perturbations, in the eikonal limit we reproduce eqs. (33) and (34) of Konoplya and Stashko (2024) and consequently confirm the correspondence. In Konoplya and Stashko (2024), the existence of non-perturbative quasinormal modes in the parameter $\xi$ was reported, which apparently cannot be reproduced by the WKB method and, therefore, do not satisfy the correspondence. However, such modes appear in the spectrum at sufficiently large values of $\xi$, while the underlying theory itself is perturbative in $\xi$. Thus, we conclude that the emergence of non-perturbative modes likely indicates a breakdown of the perturbative regime of quantum corrections, rather than a breakdown of the correspondence.
@@ V Conclusions
Recent work Zhang et al. (2024) derives and analyzes quantum-modified black hole spacetimes that address the issue of general covariance in spherically symmetric gravity by applying minimal requirements to the effective Hamiltonian constraint. While the quasinormal modes of bosonic fields have been thoroughly studied in Konoplya and Stashko (2024); Bolokhov (2024c); Malik (2024), no such analysis has been conducted for fermionic perturbations. In this paper, we complement previous studies by calculating the quasinormal modes of the massless Dirac field in the background of the two models of quantum-corrected black holes developed in Zhang et al. (2024). The data obtained through time-domain integration, with complex frequencies extracted using the Prony method, is consistent with the results from the 6th-order WKB method with Padé approximants within the range of parameters where both methods are reliable. Additionally, we derive an analytic formula for the frequencies as an expansion beyond the eikonal limit. The two models of quantum-corrected black holes can be distinguished by their quasinormal spectra.

### v2

@@ Abstract
Recently, two models of quantum-corrected Schwarzschild-like black holes were developed within Effective Quantum Gravity, and the spectra of bosonic perturbations have been analyzed in several recent studies. In this work, we investigate the quasinormal modes of a massless Dirac field perturbations around these black holes. By employing the higher-order WKB method and time-domain integration, we achieved consistency between the two approaches within their common range of validity. This concordance enables us to distinguish the two quantum-corrected black hole models from each other and from their Schwarzschild limit through their quasinormal spectra. In addition, we find approximate analytic expressions for quasinormal modes and grey-body factors in the form of expansion beyond the eikonal limit.
@@ I Introduction
Quasinormal modes (QNMs) a black hole background describe the characteristic oscillations of perturbations in the spacetime surrounding the black hole Nollert (1999); Kokkotas and Schmidt (1999); Konoplya and Zhidenko (2011). These modes are defined by complex frequencies, where the real part corresponds to the oscillation frequency, and the imaginary part is related to the damping rate due to the emission of radiation. Quasinormal modes arising from spacetime perturbations are observed through gravitational-wave interferometers Abbott et al. (2016, 2017, 2019). Combined with observations in the electromagnetic spectrum Akiyama et al. (2019a, b); Goddi et al. (2016), these detections enable the study of gravity in the strong-field regime.
@@
While the majority of publications on quasinormal modes are devoted to bosons fields, the fermionic perturbations play its role as cosmological background or astrophysical environment of a black hole. The Dirac field in this context is usually ascribed to neutrinos. Properties of neutrinos, in its turn, are intrinsically bound to a number of cosmological problems Lesgourgues and Pastor (2006); Boyarsky et al. (2009); Vagnozzi et al. (2017, 2018). The study of Dirac QNMs is crucial for understanding the stability of fermionic perturbations and provides insights into the behavior of quantum fields in curved spacetime. In the eikonal limit, the QNMs are closely related to the properties of null geodesics, particularly the orbital frequency and the Lyapunov exponent, which characterize the stability of these geodesics Cardoso et al. (2009); Konoplya and Stuchlík (2017); Konoplya (2023a); Bolokhov (2024a). However, Dirac QNMs exhibit unique features, such as spin-orbit coupling effects, that distinguish them from scalar or gravitational perturbations. These modes also play a some role in the context of black hole thermodynamics and quantum gravity, where they can be used to probe the underlying quantum structure of spacetime. Analytical and numerical methods, including the WKB approximation and time-domain integration, are commonly employed to calculate these modes, revealing how different black hole parameters, such as mass, charge, and spin, influence the fermionic QNM spectrum. As a result, quasinormal modes of Dirac fields have been extensively studied for a great number of black hole models and gravitational theories Zinhailo (2019); Bolokhov (2024b); González and Vásquez (2014); Jing and Pan (2005); Konoplya and Zhidenko (2018); Rosa and Dolan (2012); Wongjun et al. (2020); Kanti et al. (2006); Al-Badawi and Kraishan (2023); Chen and Jing (2005); Konoplya et al. (2022); Saleh et al. (2016); Chakrabarti (2009); Varghese and Kuriakose (2011); Sebastian and Kuriakose (2014); Lopez-Ortega (2007); Wu and Zhao (2004); Zhang and Jing (2005); Lopez-Ortega (2010); Wang and Gui (2010); Wang et al. (2017); Konoplya and Zhidenko (2007a); Jing (2004); Al-Badawi et al. (2023); Zhou and Zhu (2004); Konoplya et al. (2020); Lopez-Ortega (2014).
There exist many different approaches to constructing quantum-corrected black hole spacetimes. As a result, the literature on perturbations and quasinormal modes of such quantum-corrected black holes is vast. One line of research focuses on black holes obtained by accounting for spherically symmetric quantum fluctuations of the metric and matter fields in a nonperturbative manner Kazakov and Solodukhin (1994), with their quasinormal spectra analyzed in Saleh et al. (2016, 2014); Konoplya (2020). Another major direction involves higher-curvature corrected theories, which are motivated by quantum corrections in the context of string theory. Black hole perturbations in such models have been studied in various dimensions: in $2+1$ dimensions Konoplya and Zhidenko (2020); Skvortsova (2024a, b), four dimensions Konoplya and Zinhailo (2020); Zinhailo (2018); Cano et al. (2020); Wagle et al. (2022); Konoplya et al. (2019a); Matyjasek (2020); Blázquez-Salcedo et al. (2020, 2022); Konoplya (2023b), and in higher-dimensional cases Abdalla et al. (2005); Gonzalez et al. (2018); Cuyubamba et al. (2016); Yoshida and Soda (2016); Konoplya and Zhidenko (2017a); Chen et al. (2016).
A large body of recent work has also focused on quasinormal modes of black holes in the framework of asymptotically safe gravity Stashko (2024); Dubinsky (2024); Malik (2024a); Konoplya et al. (2023a); Pedrotti and Vagnozzi (2024). Additionally, various loop quantum gravity–inspired approaches, such as the Hamiltonian constraint method and holonomy corrections, have been employed to study the spectra of quantum-corrected black holes Bolokhov (2024c, d); Liu et al. (2020); Skvortsova (2024c); Konoplya and Stashko (2025); Yang et al. (2024); Gong et al. (2024); Moreira et al. (2023); Zinhailo (2024); del Corral and Olmedo (2022). Finally, the quasinormal spectrum of a quantum-corrected black hole derived from T-duality arguments has been investigated in Nicolini et al. (2019); Konoplya et al. (2023b).
@@
One the promising approaches to construction of quantum gravity is the Loop Quantum Gravity Ashtekar and Lewandowski (2004); Ashtekar (1987); Ashtekar and Lewandowski (1997), the test for which in the electromagnetic and gravitational sectors have been studied in Vagnozzi et al. (2023); Afrin et al. (2023). Recently, within the Hamiltonian constraints approach, intriguing models for spherically symmetric quantum-corrected black holes that preserve general covariance have been developed in Zhang et al. (2024). As a result, various physical properties of these black holes, such as optical phenomena, accretion and particle motion, have been studied in Liu et al. (2024a, b); Chen and Yang (2025); Shu and Huang (2025). The quasinormal modes of boson fields for these models have also been recently investigated in Konoplya and Stashko (2024); Bolokhov (2024e); Malik (2024b). In this work, we extend these studies to consider massless Dirac field perturbations, implying, first of all, propagating neutrinos in a black hole background, and calculate the corresponding quasinormal frequencies using two alternative methods: the higher-order WKB method with Padé approximants and time-domain integration. We will demonstrate that the quasinormal modes of the Dirac field exhibit a noticeable shift when the quantum parameter is introduced. In addition we obtain the compact analytic formula for quasinormal frequencies in the form of the expansion beyond the eikonal limit. Using the correspondence between the grey-body factors and quasinormal modes Konoplya and Zhidenko (2024a) this formula gives an analytic expression for the grey-body factors for neutrinos.
@@ II Black hole metric and wave-like equation
The metric of the quantum corrected black hole is given by the following line element Zhang et al. (2024)
@@ III.1 WKB Method
This expression provides the square of the QNM frequencies $\omega_{n}$ in terms of the maximum of the potential $V_{0}$, its second derivative $V_{0}^{\prime\prime}$, and the higher-order WKB corrections $\Lambda_{j}$. The sum of higher-order corrections $\sum_{j=1}^{\infty}\Lambda_{j}$ ensures the accuracy of the QNM frequency calculation, especially when going beyond the leading WKB order. The explicit forms of the corrections have been derived in Iyer and Will (1987) up to the third order, and in Konoplya (2003) and Matyjasek and Opala (2017) up to the sixth and 13th orders respectively. The WKB method has been used for finding quasinormal modes in numerous works (see, for example, Konoplya and Zhidenko (2007b); Balart et al. (2023); Bretón et al. (2017); Yu et al. (2022); Konoplya and Zhidenko (2022); Al-Badawi (2023); Xiong et al. (2022); Kokkotas et al. (2011)), so that we will not discuss it here in more detail. Here we will use the 6th order WKB method Konoplya (2003) with Padé approximants Matyjasek and Opala (2017); Konoplya et al. (2019b).
@@ III.2 Time-Domain Integration
$\Psi(N)=\Psi(W)+\Psi(E)-\Psi(S)-\Delta^{2}V(S)\frac{\Psi(W)+\Psi(E)}{8}+\mathcal{O}\left(\Delta^{4}\right),$ (9)
@@ IV Quasinormal modes
Here, we will also derive an analytic approximation for quasinormal modes using an expansion in powers of the inverse multipole number, as suggested in Konoplya and Zhidenko (2023) and applied in various studies for different black hole models Bolokhov (2024b); Malik (2024a, c, d, e) and reviewed in Bolokhov and Skvortsova (2025). Perturbations in a spherically symmetric background can be reduced to the wave-like equation with the effective potential which can be approximated in the following way:
@@
In Cardoso et al. (2009), it was demonstrated that there is a correspondence between the parameters of null geodesics (orbital frequency and Lyapunov exponent) and the real and imaginary parts of quasinormal modes in the eikonal limit. This correspondence is known to hold in the majority of cases. However, there are several exceptions and limitations, as discussed in Konoplya and Stuchlík (2017); Konoplya (2023a); Bolokhov (2024a). The eikonal limit is interesting not only because in this regime the frequencies could be found in analytic form (see, for instance, Konoplya and Abdalla (2005); Bolokhov (2024c); Konoplya (2002); Zhidenko (2004); Allahyari et al. (2019)), but also because it may bring a catastrophic instability Konoplya and Zhidenko (2017b); Takahashi (2011); Konoplya and Zhidenko (2017c); Gleiser and Dotti (2005); Dotti and Gleiser (2005) in theories with higher corrections in curvature.
@@
For the present case of Dirac field perturbations, in the eikonal limit perturbation equations do not depend on the spin of the field and we reproduce eqs. (33) and (34) of Konoplya and Stashko (2024), and, consequently, confirm the correspondence. In Konoplya and Stashko (2024), the existence of non-perturbative quasinormal modes in the parameter $\xi$ was reported, which apparently cannot be reproduced by the WKB method and, therefore, do not satisfy the correspondence. However, such modes appear in the spectrum at sufficiently large values of $\xi$, while the underlying theory itself is perturbative in $\xi$. Thus, we conclude that the emergence of non-perturbative modes likely indicates a breakdown of the perturbative regime of quantum corrections, rather than a breakdown of the correspondence.
It is worth noting that the analytically derived expressions above allow one to compute not only the fundamental quasinormal mode but also overtones, via their dependence on the parameter $K=n+\tfrac{1}{2}$. However, the WKB approximation remains sufficiently accurate only in the regime where the overtone number $n$ is smaller than the multipole number $\ell$. Consequently, it cannot capture the so-called “outburst of overtones” observed in Konoplya and Zhidenko (2024b), which refers to a strong deviation of overtone frequencies from their Schwarzschild values as $n$ increases. Although studies of integer-spin perturbations in quantum-corrected black holes Konoplya and Stashko (2024) suggest the presence of significant deviations in overtone spectra, such overtones are not typically visible in time-domain profiles. This is because the fundamental mode dominates the energy content of the ringdown signal. Indeed, we were also unable to extract overtones from time-domain profiles in the case of Dirac perturbations. However, using an alternative representation of the Dirac wave equation—one that avoids square roots—and subsequently applying the precise Leaver method Leaver (1985), would allow accurate extraction of overtone frequencies even in the regime $\ell\leq n$.
In addition to the quasinormal modes boundary conditions problem, another important spectral problem is related to the Hawking radiation of quantum corrected black holes, which has been actively studied recently in Calzà et al. (2025a, b). This is the scattering problem which allows one to find the portions of reflected and transmitted particles of the initial Hawking radiation flow and thereby to find the total Hawking evaporation rate.
The obtained here analytic expression for quasinormal modes can be effective used for finding aanalytic approximation for the grey-body factors through the correcspondence between quasinormal modes and grey-body factors Konoplya and Zhidenko (2024a). Thus taking the fundamental mode $\omega_{0\ell}$ and the first overtone $\omega_{1\ell}$ for a given multipole number $\ell$ and using eq. 4.6 of Konoplya and Zhidenko (2024a), we can find the grey-body factors as follows:
$\Gamma_{\ell}(\Omega)\equiv|T|^{2}$ $=$ $\left(1+e^{2\pi\dfrac{\Omega^{2}-Re(\omega_{0\ell})^{2}}{4Re(\omega_{0\ell})Im(\omega_{0\ell})}}\right)^{-1}+beyond~{}eikonal~{}corrections.$ (22)
Here $\Omega$ is the real frequency of the wave in the scattering problems and the beyond-eikonal corrections (which depend also on $\omega_{1\ell}$) can be found in eqs. (4.2-4.5) of Konoplya and Zhidenko (2024a). The above equation together with the obtained here values $\omega_{0}$ and $\omega_{1}$ gives the transimission coefficient for neutrinos in the background of the considered quantum corrected black holes.
@@ V Conclusions
Recent work Zhang et al. (2024) derives and analyzes quantum-modified black hole spacetimes that address the issue of general covariance in spherically symmetric gravity by applying minimal requirements to the effective Hamiltonian constraint. While the quasinormal modes of bosonic fields have been thoroughly studied in recent works Konoplya and Stashko (2024); Bolokhov (2024e); Malik (2024b), no such analysis has been conducted for fermionic perturbations. In this paper, we complement previous studies by calculating the quasinormal modes of the massless Dirac field, implying propagation of neutrinos, in the background of the two models of quantum-corrected black holes developed in Zhang et al. (2024). The data obtained through time-domain integration, with complex frequencies extracted using the Prony method, is consistent with the results from the 6th-order WKB method with Padé approximants within the range of parameters where both methods are reliable. Additionally, we derive an analytic formula for the frequencies as an expansion beyond the eikonal limit. The two models of quantum-corrected black holes can be distinguished by their quasinormal spectra. The obtained analytic formula for quasinormal modes can also be used to obtain the analytic expression for the neutrino’s grey-body factors. Our work could be extended by considering massive Dirac perturbations, which, as massive field of every spin, exhibit several qualitative differences both during the ringdown phase and in the asymptotic tail behavior Ohashi and Sakagami (2004); Konoplya et al. (2007); Bolokhov (2024c); Koyama and Tomimatsu (2001); Lütfüoğlu (2025).